Apologies for absence
There were none.
minutes of the last meeting.
These were accepted.
Report from Computing Executive Group
This item was not taken.
Reports from units.
George reported that the User Support Unit would be the initial point of contact for any problems with the VoIP telephones.
The most recent meeting to consider the planned work for the Appleton Tower basement only considered areas affecting PPLS and IS and so George was not present at that meeting.
Servers attached to UPSs in the Appleton Tower server room were shut down smoothly yesterday morning when power was lost to the whole of the basement area. The power outage lasted about an hour and a half and then it took a further half hour approximately to bring back all the services. The power outage affected much of the School and Windows machines and routing throughout many areas of the University.
A second switched rack power distribution unit has been delivered and installed in the top of the new rack in the Appleton Tower server room.
Temperature graphs derived from data from the temperature sensor in the Appleton Tower server room are now on the web.
George will be investigating ways of hooking up the VoIP phones status and the temperature sensors to the new monitoring system.
A new rack has been installed in the JCMB server room.
Stephen reported that there were still about 20 (22 according to Alastair) autobuilt packages which need to be rebuilt on the FC6 and FC6 64bit platforms and submitted to the corresponding repositories. This should be done by next Monday, October 1st.
We now have two SL5 build hosts available, buildsl5 and buildsl5_64 which can be used by anyone needing to build packages on these two platforms. SL5 has now been installed on practically all the MPU's test servers. All packages at the inf level have been done (although it currently includes autobuild versions of some of these). SL5 can now be installed via PXE or CD and works under vmware. Anyone wishing to install their own SL5 machine must use inf level headers.
Alastair is currently investigating dual path fibre channel support.
The MPU's test servers will shortly be moved to the new rack in the JCMB server room. This will create space for future expansion of server space for other units that need fibre channel connections.
Research and Teaching.
Tim reported that they have been seeing evidence of some database oddities again, causing occasional hanging and crashing. It mainly appears to be affecting computing staff rather than the administrative staff. Ken commented that he had also noticed some errors when trying to use the sql command line interface. He wondered whether it had anything to do with the recent running of the syncdbldap script in the morning. It should now be safe to remove that particular run of the script, so Tim will see whether doing this improves things at all.
The computerdesign.h header file has been introduced for machines in AT-3.09 which have been specially modified for use in the Computer Design module. The header converts the machines to using the sl5 kernel and adds support for xilinx and a few other packages. The xilinx software is too big at 7GB to distribute as an rpm and is mounted from a shared area of file space.
Tim reported that some computer design machines had been spontaneously shutting themselves down or powering themselves off. It was suggested that it was more likely that they were powering themselves off if they were overheating. Tim will look at the syslog files on the machines in question.
Work has been done on configuring the new exam preparation desktop machines to firewall them against any network traffic other than that required to manage them via LCFG. George has inspected the configuration but Tim asked that a peer review be done of the set up to look for any possible ways that access could be gained to the machines from elsewhere on the network. He will mail round details of which machine is currently set up in this way. He will also drop some of the additional but unnecessary firewall modules. Stephen mentioned that the DNS protocol was sometimes used to carry non-DNS traffic when a hacker was trying to break out of a firewalled machine.
Craig thanked Stephen and Alison for their work on getting CUPS to act as a forwarding mechanism to enable printing to an LPRng print service from OpenOffice and Firefox on FC6.
We will soon be in receipt of a Canon iRC2880 multi-function colour printer/copier leased from the Printing Services section of the Procurement Office of the University. This will be installed in Buccleuch Place and will displace a colour printer from there to the JCMB to replace a broken colour printer there.
The Sun Sunfire 280R file server wyvern is now an AFS server and is being used by the Managed Platform Unit initially. Advantage will be taken of this situation to do some investigations into fine tuning the server performance
The Dell PE 1950 mullo is being used as a test AFS linux file server and will shortly be introduced as a fully operational AFS server.
Alastair asked about the behaviour of aklog since it appeared to be taking up to 30 seconds on occasion yesterday morning to carry out its task. He will send Craig the relevant logs.
The Unit are now ready to make the Informatics web server use Cosign but are waiting for a few things to be resolved such as the configuration of firefox for SPNEGO and making sure that the Front Line Support Team are ready to handle queries from staff about the new behaviour.
Since the last Operational Meeting the User Support Unit had handled 254 new RT tickets (equivalent to about 28 per working day) and resolved 57% of them. There had been a total of 208 tickets (including both new and existing tickets) resolved over the same period.
Ken reported some figures showing resolution rates against age of recent RT tickets:
|Ticket Age D||Percentage Resolved (or rejected)|
|1 week < D < 2 weeks||71%|
|2 weeks < D < 1 month||76%|
|1 month < D < 3 months||90%|
Various bug fixes had been made to the new version of the accgen script when it was first used in earnest on the Tuesday morning of week 0 of the semester. The script now appeared to be working satisfactorily. About 5 AFS accounts out of the 270 odd accounts recently created by the create_afsuser script appeared to have not been created absolutely correctly (either missing entry from the PTS or no mirror volume created).
The total number of accounts (not including temporary ones) which have an AFS home directory is now 581 (an increase of 273 in the last two weeks). Of these accounts, 391 are for students and 190 are for non-students.
Ken briefly reported that 80% of all office machines have now been upgraded to FC5. He did not give details during the meeting but provides them below:
About 54 machines had been upgraded/installed with FC6 in the last two weeks. The following figures give the numbers of FC6 and FC5 machines (based on a simple analysis of machine profiles):
However the best estimate of office machines managed by User Support that are still at FC5 is about 50.
There was none.
The next meeting will be on October 10th in the Buccleuch Place Seminar Room and chaired by Craig Strachan.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh